Multiple-Froduct Analyss, Changes in Sales Mix, Sales to Earn Target Operating Income Basu Company produces two types of sleds for playing in the snow: basic sled and aerosled. The projected income for the coming year, segmented by product line, follows: Basic SledAerosled Total 53,000,000 2,400,000 $5,400,000 1,000,000 1,000,000 2,000,000 2,000,000 $1,400,000 3,400,000 00 550,000 1,428,000 1,222,000 $750,000 $1,972,000 198,900 $1,773,100 Total variable cost Direct fixed cost Product margin Common fixed cost Operating in The selling prices are s30 for the basic sled and $60 for the aerosled. (Round bresk-even packages and break-even units to the nearest whole unit.) Required 1. Compute the number of units of esch product that must be sold for Basu to break even. Basic 11.11 x units 23,800 X units Assume that the marketing manager changes the sales mix of the two products so that the ratio is five basic sleds to three aerosleds. Compute the number of units of each product that must be sold for Basu to break even. Round your answers to the nearest whole number. 34,8291 X units 20.098 x units of basic J. Conceptual Connection: Refer to the orginal dats. Suppose that su can increase the sales of serosleds with increased advertising. The extrs advertising would cost an additional $195,000, and some of the potential purchssers aerosleds would increase by 12,000 units, and sales of basic sleds would decrease by 5,000 units. Would Basu be better off with this strategy? f so, give the amount of increase in income. Yes v sleds would switch to serosleds. In total sales of basic eds would 125,000 7 Cheock My Work Partaly correct
Answer #1

1. Compute the number of units of each product that must be sold for Basu to break even

Basic sled

Aero Sled

Sales

$          3,000,000

$     2,400,000

Price

$                        30

$                   60

Units (C1)

                 100,000

              40,000

Variable Cost Per Unit (C2)

$                        10

$                   25

Calculate Break Even Point (BEP) for each product:

BEP = Total Fixed Cost / (Price– Variable Cost)

BEP (Basic sled) = 778,000/ (30 - 10)

                            = 38,900 Units

BEP (Aero sled) = 650,000/ (60 - 25)

                           = 18,572 Units

Number of units sold:-

Basic sled - 38,900 Units

Aero sled - 18,572 Units

Calculations:

C1 -

Units = Sales / Price

Units (Basic Sled) = 3000000 / 30

= 100,000

Units (Aero Sled) = 2,400,000 / 60

= 40,000

C2 -

Variable Cost Per unit = Variable Cost / Unit

Variable Cost Per unit (Basic Sled) = 1,000,000 / 100,000

= $ 10

Variable Cost Per unit (Aero Sled) = 1,000,000 / 40,000

= $25

2. Assume that the marketing manager changes the sale mix of the two products so that the ratio isfive basic sled to three aero sled.Compute the number of units of each product that must be sold for Basu to break even.

Basic Sled: Aero sled = 5: 3

Product

Price

Unit Variable Cost

Unit Contribution Margin (C3)

Sales Mix

Package Contribution Margin (C4)

Basic

$ 30

$ 10

$ 20

5

$ 100

Aero

$ 60

$ 25

$ 35

3

$ 105

Total

$ 205

Break Even Point (Package) = Total Fixed Cost (C5) / (Total Package Contribution Margin)

                                              = 1,636,900 / 205

                                              = 7,937

Break Even Point for Each Product = BEP (Package) * Sales Mix

BEP (Basic sled) = 7937 * 5

                            = 39,685 Units

BEP (Aero sled) = 7937 * 3

                            = 23,811 Units

Calculations:

C3 –

Unit Contribution Margin = Price – Variable Cost per unit

Unit Contribution Margin (Basic Sled) = 30 - 10

                                                               = $ 20

Unit Contribution Margin (Aero Sled) = 60 - 25

                                                               = $ 35

C4 –

Package Contribution Margin = Unit Contribution Margin * Sales Mix

Package Contribution Margin (Basic Sled) = 20 * 5

       = $ 100

Package Contribution Margin (Aero Sled) = 35 * 3

        = $ 105

C5 –

Total fixed cost = Total direct fixed cost + Common Fixed cost

                        = 1,428,000 + 1, 98,900

                        = $ 1,636,900

3. Refer to Original Data. Suppose that Basu can increase The extra advertising would cost anadditional $195,000 and some of the potential purchasers of basic sled would switch toaerosleds. In total, sales of aerosled would increaseby 12,000 units, and sales of basic sledswould decrease by 5000 units. Would Basu be btter with this strategy? If so, give the amount of increase in income.

Before Advertising:

Basic

Aero

Total

Units Sold

100,000

40,000

Per Unit CM

$ 20

$ 35

Total (C6)

$ 2,000,000

$ 1,400,000

Direct Fixed Cost

$ 778,000

$ 650,000

Product Margin (C7)

$ 1,222,000

$ 750,000

1,972,000

Common Fixed Cost

$ 198,900

Operating Income (Increased By) (C8)

$ 1,773,100

After Advertising:

Basic

Aero

Total

Units Sold

95,000

52,000

Per Unit CM

$ 20

$ 35

Total (C6)

$ 1,900,000

$ 1,820,000

Direct Fixed Cost

$ 778,000

$ 650,000

Product Margin (C7)

$ 1,122,000

$ 1,170,000

2,292,000

Common Fixed Cost

$ 198,900

Advertising Cost

195,000

Operating Income (Increased By) (C8)

$ 1,898,100

The operating income has increased when the sales of aero sled was increased by 12,000 units and the additional advertising cost was incurred.

So we can say basu can go-ahead with this strategy.

The amount of increase in income of Basu is $ 125,000 (C9)

Calculations:

C6 –

Total = Unit Sold * Per Unit CM

Before Advertising –

Total (Basic Sled) = 100000 * 20

= 2,000,000

Total (Aero Sled) = 40,000 * 35

            = 1,400,000

After Advertising –

Total (Basic Sled) = 95000 * 20

            =1,900,000

Total (Aero Sled) = 52000 * 35

            = 1,820,000

C7 –

Product Margin = Total – Direct Fixed Cost

Before Advertising -

Product Margin (Basic Sled) = 2,000,000 – 7, 78,000

= $ 1,222,000

Product Margin (Aero Sled) = 1,400,000 – 650,000

            = $ 750,000

After Advertising -

Product Margin (Basic Sled) = 1,900,000 – 778,000

            =$ 1,122,000

Product Margin (Aero Sled) = 1,820,000 = 650,000

            = $ 1,170,000

C8 –

Operating Income (Increased By) –

Before Advertising = Total product Margin – Common Fixed Cost

            =1,972,000 - 198,900

= $ 1,773,100

After Advertising = Total product Margin – Common Fixed Cost – Advertising Cost

            = 2,292,000 - 198,900 - 195,000

            = $ 1,898,100

Calculation:

C9 –

Amount of increase in income.

= Operating Income after Advertising - Operating Income before Advertising

= 1,898,100 - 1,773,100

= $ 125,000